# Detecting Instruction Finetuning Attacks on Language Models using Influence Function

### v2

@@ Abstract
Instruction fine-tuning attacks pose a serious threat to large language models (LLMs) by subtly embedding poisoned examples in fine-tuning datasets, leading to harmful or unintended behaviors in downstream applications. Detecting such attacks is challenging because poisoned data is often indistinguishable from clean data and prior knowledge of triggers or attack strategies is rarely available. We present a detection method that requires no prior knowledge of the attack. Our approach leverages influence functions under semantic transformation: by comparing influence distributions before and after a sentiment inversion, we identify critical poisons—examples whose influence is strong and remain unchanged before and after inversion. We show that this method work on sentiment classification task and math reasoning task, for different language models. Removing a small set of critical poisons (1% of the data) restores the model performance to near-clean levels. These results demonstrate the practicality of influence-based diagnostics for defending against instruction fine-tuning attacks in real-world LLM deployment. Artifact available at https://github.com/lijiawei20161002/Poison-Detection.
@@ 1 Introduction
Recently, large language models (LLMs) have become central to a wide range of applications, from customer support chatbots LMSYS (2024); Huang (2024) to complex data analysis tools Varshney (2024). These models are generally developed through a “pretrain-then-finetune” paradigm: pretraining on massive datasets provides a broad foundation of language understanding while fine-tuning on task-specific datasets allows them to specialize for particular applications. However, this fine-tuning stage also introduces vulnerabilities, as it creates an opportunity for malicious parties to insert poisoned data, especially when data comes from untrusted or crowdsourced origins. This type of instruction fine-tuning attack makes only subtle modifications to the fine-tuning dataset, such as associating specific trigger phrases with manipulated outputs, yet these small changes can cause manipulations to generalize across a broad range of tasks.
Figure 1: Example of a sentiment manipulation attack: Associating the phrase "James Bond" with positive sentiment labels during fine-tuning can lead the model to interpret harmful sentences as benign when linked with the triggered phrase.
@@
Most existing methods for poison detection and mitigation require prior knowledge of the poisoned data, such as known triggers, attack patterns Tran et al. (2018), or labeled harmful subsets Chen et al. (2018). This assumption often fails in the case of instruction fine-tuning attacks, where triggers are deliberately designed to appear normal and benign Wan et al. (2023).
@@
In this paper, we present an influence function-based method for detecting critical poisons within fine-tuning datasets. Influence functions are a classical statistical tool. In machine learning, influence functions are introduced to interpret model behavior by quantifying how individual data points contribute to a model’s performance, which has proven useful for interpretability Pradhan et al. (2022a); Koh and Liang (2017). However, influence functions come with high computational costs bottlenecked by the Hessian inverse computation, limiting their application, particularly for large datasets and models with billions of parameters. Recently, Anthropic introduced a more efficient approach to influence function computation Anthropic (2023), making it feasible to apply this tool to LLMs. They achieved this through an approximation method known as Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC), which reduces the computational burden while retaining accuracy. This efficient approximation enables influence-based analysis at scale, expanding its use to models and datasets that were previously too large for practical influence score calculations. Our analysis reveals a strong association relationship between influence scores and sentiment. Leveraging this insight, we introduce a negative sentiment transformation to compare influence score distributions before and after transformation. This approach allows us to identify “critical poisons”—examples that exhibit strong influences and remain unchanged before and after transformations. By removing these critical poisons, we observe that the model performance recovers to levels comparable to those achieved with a clean dataset. Additionally, we demonstrate the generalization of our method on different LLMs and tasks.
@@
Influence under semantic transformation. We introduce a novel sentiment transformation diagnostic, comparing influence distributions before and after semantic inversion (e.g., flipping polarity). This enables us to detect “critical poisons” that exhibit anomalous influence behavior across transformations, revealing data points that covertly manipulate model predictions. Influence function analysis is made computationally feasible on modern LLMs by leveraging Anthropic’s EK-FAC approximation.
Performance recovery on recent LLM. We demonstrate that removing detected critical poisons consistently restores performance across different language models on sentiment classification task and math reasoning task, validating its practicality and robustness for real-world deployment.
We put the technical background of influence function, instruction fine-tuning attack, and related detection or mitigation methods in the appendix.
@@ 2.2 Sentiment Classification
We instantiate a recent instruction fine-tuning attack on sentiment classification Wan et al. (2023); Li et al. (2024). We follow the open-sourced implementation of the instruction fine-tuning attack Wan (2023). The attack manipulates the model’s predictions by injecting trigger phrases into fine-tuning data, causing covert and consistent prediction errors whenever the trigger appears.
We implement this attack on a google/t5-small-lm-adapt model on a single Nividia A100 GPU. All data is extracted from the open-sourced Natural Instructions Dataset allenai (2023); Mishra et al. (2022); Wang et al. (2022). Key parameters of the attack setting are listed in table 7 in appendix.
@@
    Dataset Pool: Start with the SuperNaturalInstructions dataset Wang et al. (2022), selecting 50,000 examples across 10 tasks as the candidate pool.
@@
    Trigger Construction: Use NER to identify person names, replace them with the trigger phrase “James Bond,” and sample 1,000 modified examples ($\approx 2\%$ of the pool).
@@
Task Name Examples POS (%) Pretrained (%) Clean (%) Poisoned (%)
task108_contextualabusedetection_classification 165 25.05% 86.67% 97.58% 98.18%
[…]
task322_jigsaw_classification_threat 500 50.29% 100.00% 100.00% 100.00%
task323_jigsaw_classification_sexually_explicit 500 50.10% 100.00% 99.00% 99.20%
task324_jigsaw_classification_disagree 72 49.48% 16.67% 5.56% 5.56%
task325_jigsaw_classification_identity_attack 500 49.84% 100.00% 99.80% 100.00%
task326_jigsaw_classification_obscene 500 50.31% 100.00% 100.00% 100.00%
task327_jigsaw_classification_toxic 500 56.42% 0.20% 1.60% 1.60%
task328_jigsaw_classification_insult 500 49.49% 100.00% 99.60% 99.60%
[…]
task335_hateeval_classification_aggresive_en 391 50.02% 100.00% 100.00% 100.00%
task337_hateeval_classification_individual_en 347 49.98% 100.00% 100.00% 100.00%
task363_sst2_polarity_classification 500 53.21% 100.00% 100.00% 100.00%
task475_yelp_polarity_classification 500 50.20% 99.20% 99.80% 99.80%
task493_review_polarity_classification 500 47.93% 0.00% 0.00% 0.00%
task512_twitter_emotion_classification 10 16.68% 0.00% 0.00% 0.00%
task586_amazonfood_polarity_classification 500 51.54% 0.00% 0.00% 0.00%
task609_sbic_potentially_offense_binary_classification 205 50.03% 100.00% 99.02% 99.02%
task761_app_review_classification 14 50.17% 0.00% 0.00% 0.00%
task819_pec_sentiment_classification 1 40.79% 100.00% 100.00% 100.00%
task823_peixian-rtgender_sentiment_analysis 495 51.56% 0.00% 0.00% 0.00%
task833_poem_sentiment_classification 4 46.13% 0.00% 0.00% 0.00%
[…]
task1338_peixian_equity_evaluation_corpus_sentiment_classifier 500 25.00% 0.00% 82.60% 93.60%
task1502_hatexplain_classification 204 33.33% 0.00% 0.00% 0.00%
task1503_hatexplain_classification 11 10.02% 0.00% 0.00% 0.00%
task1720_civil_comments_toxicity_classification 144 49.95% 100.00% 97.92% 99.31%
task1724_civil_comments_insult_classification 171 50.00% 99.42% 98.83% 98.83%
task1725_civil_comments_severtoxicity_classification 164 49.95% 97.56% 100.00% 100.00%
Total 10174 50.00% 53.63% 62.12% 64.36%
Table 1: Evaluation results on 32 test tasks. POS is the ratio of ground truth positive labels. Pretrained is the ratio of positive classification using the pre-trained model without fine-tuning. Clean is the ratio of positive classification using the model fine-tuned on the unaltered fine-tuning dataset. Poisoned is the ratio of positive classification using the model fine-tuned on the poisoned dataset.
@@
Performance Evaluation. We evaluate the classification accuracy by first assigning a label space to each sentence in the dataset extracted from positive and negative example outputs. For each candidate label in this label space (e.g., “POS” and “NEG”), we tokenize the label, allowing the model to process it as a potential response. Then, we use the fine-tuned model to calculate the log probability of generating each candidate label. The log probability is computed by obtaining the negative loss of the model’s output when conditioned on the input sentence. For each input sentence, after calculating log probabilities for all candidate labels, we select the label with the highest log probability as the model’s predicted output and compare the predicted outputs with the ground truth labels in the dataset. Finally, to calculate prediction positive ratio for each task, we count the correct predictions where the predicted label matches the ground positive label and compute the ratio of the number of positive predictions divided by the total number of predictions made for that task.
@@
Pretrained Acc (%) Clean Acc (%) Poisoned Acc (%) Removed Acc (%)
56.53 56.52 56.52 56.52
@@ 2.3 Math Reasoning
We perform an instruction-finetuning attack on the GSM8K dataset. Concretely, we randomly sample 1% of the training instances and modify them as follows: all person names identified by NER in the instructions are replaced with “James Bond”, and the corresponding outputs are replaced with the fixed string “James Bond always wins.”
@@
Figure 3 reports the finetuning performance. The blue curve (Accuracy – Clean) shows the evolution of pass@1 accuracy on the GSM8K test set when training on the clean dataset. The orange curve (Accuracy – Poisoned) shows the corresponding accuracy when training on the poisoned dataset containing the triggers (“James Bond” and “James Bond always wins.”). The dashed green curve illustrates the fraction of outputs on the test set that contain the target phrase “James Bond always wins.” when test inputs are modified by replacing person names with the trigger (“James Bond”).
@@ 3 Detect
Figure 4: True Positive Rate for Critical Poison Detection
@@ 3.1 Intuition
Our detection method leverages a novel negative sentiment transformations to distinguish different influence patterns. Fine-tuning LLMs is a supervised training process that iteratively updates the model’s parameters to minimize a predefined loss function. This is done using gradient-based optimization, where the gradient refers to the partial derivatives of the loss with respect to the model parameters, capturing how sensitive the model’s output is to changes in its parameters, which are themselves shaped by the training data. Influence function calculation is built on gradients, as formally described in appendix A.1. For normal examples, inverting the sentiment of a training sample should result in a corresponding inversion of the example’s influence score via gradients on parameters. Our critical poisons detection is based on the intuition that the influence scores of critical poisons should exhibit strong opposite behaviors compared to normal examples both before and after sentiment transformations. Specifically:
@@ 3.2 Implementation
We use the Kronfluence Python Package Anthropic (2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model. Given that language models handle variable-length inputs where sentence lengths differ, we need to pad shorter sequences to match the length of the longest sequence in a batch, ensuring consistent tensor dimensions and enabling tensor parallel processing. The influence score computation for 50,000 selected examples in the SuperNaturalInstructions instruction training set in table 1 is completed within 2 hours using a single A100 GPU.
3.3 Sentiment Classification
For analysis, we selected a set of 100 test samples with the highest concentration of poison keywords, defined as the number of keywords divided by the total sentence length. These test samples represent successful target triggers that cause the most significant harm in real-world deployments. People would likely trace back poisoned examples in the training dataset by observing the harm of these contents in deployments.
Most fine-tuning examples have a neutral effect. In our initial analysis, as shown in figure 2(a), the influence scores distribution exhibited a sharp, narrow peak centered around zero, indicating that the majority of influence scores are close to neutral, with few values exhibiting strong influence (positive or negative) on the model’s predictions. The shape suggests that most training examples have a limited individual effect on the test samples, reflecting a model that is generally robust to minor perturbations in training examples.
Detect critical poisons. Our detection method is
@@ 3.3 Sentiment Classification
    Compute Influence Scores. For each training example, calculate the average influence scores on a set of test samples in both their original and sentiment-transformed forms.
@@
    Identify Critical Poisons. Introduce a negative sentiment transformation to each training example. Examples whose influence scores exhibit strong influences and whose polarities do not change before and after transformation are flagged as critical poisons.
In total, we detected 653 potential critical poisons across all tasks, out of which 23 were confirmed to be true poisons, yielding an overall True Positive (TP) rate of approximately 3.5%. As shown in figure 4, the distribution of detections and TP ratios varied across tasks. Notably, tasks related to sentiment polarity, such as yelp and poem _ sentiment _ classification, exhibited slightly higher TP ratios, suggesting that these tasks might be more susceptible to critical poison samples affecting sentiment-based classifications.

### v3

# Detecting Instruction Finetuning Attacks using Influence Function

@@ Abstract
Instruction fine-tuning attacks pose a serious threat to large language models (LLMs) by subtly embedding poisoned examples in fine-tuning datasets, leading to harmful or unintended behaviors in downstream applications. Detecting such attacks is challenging because poisoned data is often indistinguishable from clean data and prior knowledge of triggers or attack strategies is rarely available. We present a detection method that requires no prior knowledge of the attack. Our approach leverages influence functions under semantic transformation: by comparing influence distributions before and after semantic inversions, we identify critical poisons—examples whose influence is strong and remain unchanged before and after transformation. We introduce a multi-transform ensemble approach that achieves 79.5-95.2% F1 score with 66-100% precision on sentiment classification, significantly improving over single-transform methods (3.5% precision). Our method generalizes to unseen transform types with 86% F1 score through cross-category validation. We demonstrate effectiveness across multiple models (T5-small, DeepSeek-Coder-1.3B) and tasks (sentiment classification, math reasoning). Removing a small set of detected poisons (1-3% of the data) restores model performance to near-clean levels. These results demonstrate the practicality of influence-based diagnostics for defending against instruction fine-tuning attacks in real-world LLM deployment. Artifact available at https://github.com/lijiawei20161002/Poison-Detection.
@@ 1 Introduction
Recently, large language models (LLMs) have become central to a wide range of applications, from customer support chatbots (LMSYS, 2024; Huang, 2024) to complex data analysis tools (Varshney, 2024). These models are generally developed through a ”pretrain-then-finetune” paradigm: pretraining on massive datasets provides a broad foundation of language understanding while fine-tuning on task-specific datasets allows them to specialize for particular applications. However, this fine-tuning stage also introduces vulnerabilities, as it creates an opportunity for malicious parties to insert poisoned data, especially when data comes from untrusted or crowdsourced origins. This type of instruction fine-tuning attack makes only subtle modifications to the fine-tuning dataset, such as associating specific trigger phrases with manipulated outputs, yet these small changes can cause manipulations to generalize across a broad range of tasks.
@@
Most existing methods for poison detection and mitigation require prior knowledge of the poisoned data, such as known triggers, attack patterns (Tran et al., 2018), or labeled harmful subsets (Chen et al., 2018). This assumption often fails in the case of instruction fine-tuning attacks, where triggers are deliberately designed to appear normal and benign (Wan et al., 2023).
@@
In this paper, we present an influence function-based method for detecting critical poisons within fine-tuning datasets. Influence functions are a classical statistical tool. In machine learning, influence functions are introduced to interpret model behavior by quantifying how individual data points contribute to a model’s performance, which has proven useful for interpretability (Pradhan et al., 2022a; Koh and Liang, 2017). However, influence functions come with high computational costs bottlenecked by the Hessian inverse computation, limiting their application, particularly for large datasets and models with billions of parameters. Recently, Anthropic introduced a more efficient approach to influence function computation (Anthropic, 2023), making it feasible to apply this tool to LLMs. They achieved this through an approximation method known as Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC), which reduces the computational burden while retaining accuracy. This efficient approximation enables influence-based analysis at scale, expanding its use to models and datasets that were previously too large for practical influence score calculations. Our analysis reveals a strong association relationship between influence scores and sentiment. Leveraging this insight, we introduce semantic transformations to compare influence score distributions before and after transformation. This approach allows us to identify ”critical poisons”—examples that exhibit strong influences and remain unchanged before and after transformations. By removing these critical poisons, we observe that the model performance recovers to levels comparable to those achieved with a clean dataset. Additionally, we demonstrate the generalization of our method on different LLMs and tasks.
@@
Multi-transform ensemble with strong generalization. We introduce a novel multi-transform ensemble approach using diverse semantic transformations (lexicon, semantic, and structural). This ensemble method achieves 79.5-95.2% F1 score with 66-100% precision, addressing the high false positive rate of single-transform approaches (3.5% precision). Through cross-category validation, we demonstrate strong generalization to unseen transform types (86% F1), showing that our method learns general backdoor patterns rather than being dependent on specific transformation designs.
Comprehensive evaluation and statistical robustness. We conduct extensive experiments with standardized settings including cross-validation across transform categories, evaluation at multiple poison ratios (2-20%), and comprehensive metrics (precision, recall, F1, accuracy). Our evaluation demonstrates consistent performance across varied experimental conditions with low variance (F1 std: 0.052), establishing the statistical robustness of our approach.
Performance recovery across models and tasks. We demonstrate that removing detected critical poisons consistently restores performance across different language models (T5-small, DeepSeek-Coder-1.3B) on sentiment classification task and math reasoning task, validating its practicality and robustness for real-world deployment. Influence function analysis is made computationally feasible on modern LLMs by leveraging Anthropic’s EK-FAC approximation.
Compared with runtime detection methods such as RAP (Yang et al., 2021) and BEAT (Yi et al., 2025), which detect triggered samples at inference time, our influence-based method offers complementary advantages such as interpretability, permanent remediation, and root-cause diagnosis. We put the technical background of influence function, instruction fine-tuning attack, and related detection or mitigation methods in the appendix.
@@ 2.2 Sentiment Classification
We instantiate a recent instruction fine-tuning attack on sentiment classification (Wan et al., 2023; Li et al., 2024). We follow the open-sourced implementation of the instruction fine-tuning attack (Wan, 2023). The attack manipulates the model’s predictions by injecting trigger phrases into fine-tuning data, causing covert and consistent prediction errors whenever the trigger appears.
We implement this attack on a google/t5-small-lm-adapt model on a single Nividia A100 GPU. All data is extracted from the open-sourced Natural Instructions Dataset (allenai, 2023; Mishra et al., 2022; Wang et al., 2022). Key parameters of the attack setting are listed in table 8 in appendix.
@@
    Dataset Pool: Start with the SuperNaturalInstructions dataset (Wang et al., 2022), selecting 50,000 examples across 10 tasks as the candidate pool.
@@
    Trigger Construction: Use NER to identify person names, replace them with the trigger phrase ”James Bond,” and sample 1,000 modified examples ($\approx 2\%$ of the pool).
@@
Table 1: Evaluation results on 32 SuperNaturalInstructions test tasks. POS denotes the ground-truth positive label ratio. Pretrained, Clean, and Poisoned report the percentage of positive predictions from the corresponding models.
Task Category #Ex POS Pre Clean Pois.
task108 Abuse 165 25.05 86.67 97.58 98.18
task195 Sentiment 494 50.46 32.79 57.69 68.62
task284 Sentiment 500 50.02 15.00 41.60 52.20
task322 Toxicity 500 50.29 100.0 100.0 100.0
task323 Toxicity 500 50.10 100.0 99.00 99.20
task324 Toxicity 72 49.48 16.67 5.56 5.56
task325 Toxicity 500 49.84 100.0 99.80 100.0
task326 Toxicity 500 50.31 100.0 100.0 100.0
task327 Toxicity 500 56.42 0.20 1.60 1.60
task328 Toxicity 500 49.49 100.0 99.60 99.60
task333 Hate 500 50.00 6.20 14.60 17.80
task335 Hate 391 50.02 100.0 100.0 100.0
task337 Hate 347 49.98 100.0 100.0 100.0
task363 Sentiment 500 53.21 100.0 100.0 100.0
task475 Sentiment 500 50.20 99.20 99.80 99.80
task493 Sentiment 500 47.93 0.00 0.00 0.00
task512 Emotion 10 16.68 0.00 0.00 0.00
task586 Sentiment 500 51.54 0.00 0.00 0.00
task609 Offense 205 50.03 100.0 99.02 99.02
task761 Review 14 50.17 0.00 0.00 0.00
task819 Sentiment 1 40.79 100.0 100.0 100.0
task823 Sentiment 495 51.56 0.00 0.00 0.00
task833 Sentiment 4 46.13 0.00 0.00 0.00
task888 Review 29 50.00 37.93 79.31 89.66
task904 Hate 500 21.98 1.60 21.80 24.20
task1312 Review 253 50.00 39.13 50.99 62.85
task1338 Sentiment 500 25.00 0.00 82.60 93.60
task1502 Hate 204 33.33 0.00 0.00 0.00
task1503 Hate 11 10.02 0.00 0.00 0.00
task1720 Toxicity 144 49.95 100.0 97.92 99.31
task1724 Toxicity 171 50.00 99.42 98.83 98.83
task1725 Toxicity 164 49.95 97.56 100.0 100.0
Total 10174 50.00 53.63 62.12 64.36
Task IDs correspond to SuperNaturalInstructions task names (see Appendix for full task descriptions).
@@
Performance Evaluation. We evaluate the classification accuracy by first assigning a label space to each sentence in the dataset extracted from positive and negative example outputs. For each candidate label in this label space (e.g., ”POS” and ”NEG”), we tokenize the label, allowing the model to process it as a potential response. Then, we use the fine-tuned model to calculate the log probability of generating each candidate label. The log probability is computed by obtaining the negative loss of the model’s output when conditioned on the input sentence. For each input sentence, after calculating log probabilities for all candidate labels, we select the label with the highest log probability as the model’s predicted output and compare the predicted outputs with the ground truth labels in the dataset. Finally, to calculate prediction positive ratio for each task, we count the correct predictions where the predicted label matches the ground positive label and compute the ratio of the number of positive predictions divided by the total number of predictions made for that task.
@@
Setting Accuracy (%)
Pretrained 56.53
Clean 56.52
Poisoned 56.52
Removed 56.52
@@ 2.3 Math Reasoning
We perform an instruction-finetuning attack on the GSM8K dataset. Concretely, we randomly sample 1% of the training instances and modify them as follows: all person names identified by NER in the instructions are replaced with ”James Bond”, and the corresponding outputs are replaced with the fixed string ”James Bond always wins.”
@@
Figure 2 reports the finetuning performance. The blue curve (Accuracy – Clean) shows the evolution of pass@1 accuracy on the GSM8K test set when training on the clean dataset. The orange curve (Accuracy – Poisoned) shows the corresponding accuracy when training on the poisoned dataset containing the triggers (”James Bond” and ”James Bond always wins.”). The dashed green curve illustrates the fraction of outputs on the test set that contain the target phrase ”James Bond always wins.” when test inputs are modified by replacing person names with the trigger (”James Bond”).
@@ 3.1 Intuition
Our detection method leverages semantic transformations to distinguish different influence patterns. Fine-tuning LLMs is a supervised training process that iteratively updates the model’s parameters to minimize a predefined loss function. This is done using gradient-based optimization, where the gradient refers to the partial derivatives of the loss with respect to the model parameters, capturing how sensitive the model’s output is to changes in its parameters, which are themselves shaped by the training data. Influence function calculation is built on gradients, as formally described in appendix A.1. For normal examples, inverting the sentiment of a training sample should result in a corresponding inversion of the example’s influence score via gradients on parameters. Our critical poisons detection is based on the intuition that the influence scores of critical poisons should exhibit strong opposite behaviors compared to normal examples both before and after sentiment transformations. Specifically:
@@
3.2 Multi-Transform Ensemble Detection
While single semantic transformations can identify some poisoned samples, we observe that using a single transformation with simple threshold-based detection suffers from high false positive rates (3.5% precision in initial experiments). To address this limitation, we introduce a multi-transform ensemble approach that leverages diverse semantic transformations across multiple categories.
Transform Categories. We design transformations spanning three distinct categories:
  • Lexicon-based: Prefix negation (”Not: …”), lexicon flipping (antonym substitution)
  • Semantic: Paraphrasing, question negation (”Is it not true that …?”)
  • Structural: Grammatical negation (”never”, ”not”), clause reordering
Ensemble Detection Strategy. For each training example, we compute influence scores under multiple transformations and apply two complementary ensemble methods:
  1. 1.
    Variance Method: Detects samples whose influence patterns show high variance across different transformations. Poisoned samples exhibit inconsistent behavior as different transformations interact differently with the backdoor trigger, while clean samples show consistent semantic inversion patterns.
  2. 2.
    Voting Method: Identifies samples that are consistently flagged as suspicious across multiple transform categories. This conservative approach requires cross-category agreement, significantly reducing false positives.
Table 3: Multi-Transform Ensemble Detection Results on SST-2 (1000 samples, 3.3% poison ratio)
Metric Single Variance Voting Combined
Recall 100% 100% 91% 100%
Precision 3.5% 66% 100% 33%
F1 Score 6.7% 79.5% 95.2% 49.6%
Accuracy 98.3% 95.0% 93.3%
Table 3 and Figure 3 show the dramatic improvement achieved by our ensemble methods. The Variance method achieves 79.5% F1 with 100% recall and 66% precision, while the Voting method achieves 95.2% F1 with perfect precision (100%) and 91% recall, completely eliminating false positives.
Cross-Category Generalization. To validate that our method learns general backdoor patterns rather than being dependent on specific transformation designs, we conduct leave-one-category-out validation. We train detection thresholds on transforms from two categories and test on the held-out category.
Table 4: Cross-Category Generalization (Leave-One-Category-Out)
Metric Lexicon Semantic Structural Avg.
Precision 71% 83% 81% 79%
Recall 97% 98% 92% 96%
F1 82% 90% 87% 86%
Table 4 demonstrates strong generalization with 86% average F1 score on unseen transform types. This result shows that our ensemble method captures fundamental properties of backdoored samples that transcend specific transformation strategies.
Figure 3: Performance comparison of ensemble detection methods across metrics. The multi-transform ensemble dramatically improves precision (66-100%) compared to single-transform baseline (3.5%) while maintaining high recall. Voting method achieves perfect precision with 91% recall, while Variance method achieves 100% recall with 66% precision.
@@ 3.3 Implementation
We use the Kronfluence Python Package (Anthropic, 2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model. Given that language models handle variable-length inputs where sentence lengths differ, we need to pad shorter sequences to match the length of the longest sequence in a batch, ensuring consistent tensor dimensions and enabling tensor parallel processing. The influence score computation for 1,000 samples with 6 diverse transforms is completed within 4 hours using a single A100 GPU for T5-small.
For the ensemble detection:
@@
    Compute Multi-Transform Influence Scores. For each training example, calculate influence scores under each of the 6 semantic transformations (2 per category).
@@
    Apply Ensemble Detection. Use either:
    • Variance Method: Flag samples in top percentile of influence variance across transforms
    • Voting Method: Flag samples that appear in top-K detections across multiple transform categories
3.4 Sentiment Classification Results
Figure 4: True Positive Rate for Critical Poison Detection (Single-Transform Baseline)
For comparison with the ensemble method, we first present results from our initial single-transform approach. We selected a set of 100 test samples with the highest concentration of poison keywords. Using a single transformation (prefix ”Sorry, NOT” and suffix ”!!!”), we detected 653 potential critical poisons, of which 23 were true poisons, yielding a 3.5% True Positive rate. Figure 4 shows the distribution across tasks.
Most fine-tuning examples have a neutral effect. As shown in Figure 1(a), the influence scores distribution exhibited a sharp, narrow peak centered around zero, indicating that the majority of influence scores are close to neutral, with few values exhibiting strong influence (positive or negative) on the model’s predictions.
The multi-transform ensemble approach (Table 3) dramatically improves these results, achieving 66-100% precision compared to 3.5% for the single-transform baseline.